OpenStudy (anonymous):

A bag contains 4 white, 3 blue, and 5 red marbles. Find the probability of choosing a red marble, then a white marble if the marbles are replaced A. 1/12 B.5/36 C. 5/6 D. 5/12

OpenStudy (anonymous):

Whoever gets this right gets fan and medal

OpenStudy (anonymous):

There are 12 total marbles. Of these 12, you have 5 reds and 4 whites. This means the probability of drawing a red marble is \(\dfrac{5}{12}\), and the probability of drawing a white marble is \(\dfrac{4}{12}=\dfrac{1}{3}\). Since the marbles are being replaced, these probabilities are independent, so \[P(\text{red then white})=P(\text{red})\times P(\text{white})\]